# X-SAM: From Segment Anything to Any Segmentation

### v1

@@ 4.3 Abaltions
Mixed Fine-tuning. We ablate the impact of mixed fine-tuning on X-SAM’s performance. As shown in Tab. 6, mixed fine-tuning improves performance on out-of-domain COCO benchmarks, demonstrating X-SAM’s robust segmentation capabilities—for example, a 6.0% AP increase on A150-OV and an 8.9% gIoU increase on Reason-Val. However, it results in a 0.8% PQ decrease on COCO-Pan due to the challenge of balancing performance in multisource training.
Dual Encoders. We ablate the design of the dual encoders in X-SAM. As shown in Tab. 7, dual encoders with either a SAM or Swin encoder benefit VGD segmentation, achieving 7.2% AP and 7.9% AP on COCO-VGD, respectively. Moreover, dual encoders with a SAM encoder consistently improve performance on GCG-Val and A150-OV, while the Swin encoder, which lacks robust segmentation capabilities, provides only a small improvement on A150-OV and even has a negative impact on GCG-Val.
Multi-stage Training. We ablate the impact of the multi-stage training strategy. As shown in Tab. 8, the S1 segmentor fine-tuning phase boosts the segmentation capability, producing a notable improvement of 9.3% PQ in COCO-Pan and 1.5% AP in the A150-OV datasets. Meanwhile, the S2 alignment pre-training phase enhances image understanding capabilities, contributing an additional 2.1% Accuracy on Conv.-MMB. By integrating these stages, X-SAM demonstrates robust advances in image segmentation and comprehension, establishing its effectiveness in addressing complex visual tasks.
Segmentor Architecture. We ablate the impact of segmentor architecture by performing segmentor fine-tuning for 12 epochs. As shown in Tab. 9, M2F decoder brings a large improvement with 9.2% PQ as the effective design of M2F. The convolution connector performs better than the MLP connector, as the convolution spatial-awareness benefits segmentation, and multi-scale further improves the performance(10.7% PQ) with more diverse scale features.
Tab. 8: Ablation on Multi-Stage(M-Stage) Training. S1: Stage 1, S2: Stage 2, S3: Stage 3, Conv.: conversation.
@@
Tab. 9: Ablation on Segmentor Architecture. Conn.: connector, M-Scale: multi-scale, Con.: convolution, M2F: Mask2Former.
@@
More ablation results can be found in the Appendix A.6.
5 Conclusion
@@
Tab. 10: The Datasets Specification. The datasets used in multi-stage training. # Sample is the total number of samples in this task.
Task Datasets # Sample
Stage 1: Segmentor Fine-tuning
Generic Segmentation COCO Panoptic (Kirillov et al. 2019) (118K) 118K
Stage 2: Alignment Pre-training
Image Conversation LLaVA  (Liu et al. 2023b) (558k) 558k
Stage 3: Mixed Fine-tuning
Image Conversation LLaVA-1.5 (Liu et al. 2023b) (665K) 665K
Generic Segmentation COCO Panoptic (Kirillov et al. 2019) (118K) 118K
VGD Segmentation COCO-VGD (117K) 117K
Referring Segmentation RefCOCO(17K), RefCOCO+(17K), RefCOCOg(22k) 301K
GCG Segmentation Grand-f (Rasheed et al. 2024) (1K), RefCOCOg (19K), PSG (28K), Flickr (148K) 195K
Reasoning Segmentation LISA ReasonSeg (Lai et al. 2024) (0.2K) 0.2K
Tab. 11: The Hyper-parameters in Multi-stage Training of X-SAM.
Item Stage 1 Stage 2 Stage 3
Segmentor Fine-tuning Alignment Pre-training Mixed Fine-tuning
batch size 64 256 64
training epochs 36 1 1
lr of dual encoders 1e-5 - 4e-6
lr of other modules 1e-4 1e-3 4e-5
optimizer AdamW (Loshchilov et al. 2017)
optimizer momentum $\beta_{1}=0.9,\beta_{2}=0.999$
weight decay 0.05 0 0.05
warmup ratio 0.03 0.03 0.03
clip max norm 0.01 1 1
image augmentation random scale[0.1, 2.0] - -
Fig. 4: The Multi-stage Training of X-SAM. X-SAM performs a multi-stage training process, including segmentor fine-tuning, alignment pre-training, and mixed fine-tuning. Segmentor fine-tuning: train the segmentor on the segmentation datasets to obtain a generalized segmentor. Alignment pre-training: train the dual projectors to align the vision features and the LLM features. Mixed fine-tuning: fine-tune the dual projectors, the segmentation decoder, and the LLM on the mixed datasets.
@@ Appendix A Technical Appendices and Supplementary Material
A.1 More Dataset Details
Training Datasets. In Tab. 10, we show the datasets used in our multi-stage training. For the segmentor fine-tuning stage, we fine-tune the segmentor on generic segmentation datasets. For the alignment pre-training stage, we pre-train the dual projectors on the alignment LLaVA 558K (Liu et al. 2023b) dataset. For the mixed fine-tuning stage, we fine-tune the whole model on mixed datasets, including both segmentation and conversation datasets. There are six types of datasets in total, including one image-level dataset and five segmentation datasets.
Building COCO-VGD Dataset. The COCO-VGD dataset is built on the images and annotations of the COCO2017 instance segmentation dataset, which provides instance-level segmentation masks for each object in an image. We automatically generate four types of visual prompts: point, scribe, box, and mask, for each instance in the image, following  (Zhang et al. 2024d). We randomly sample one type of visual prompt for each category as the visually grounded prompt during training and evaluation.
Dataset Balance Resampling. As shown in Tab. 10, the dataset sizes vary, ranging from 0.2K to 665K. The data ratio among the datasets is crucial to the model’s performance during mixed fine-tuning. To balance the different dataset sizes, we propose a dataset balance resampling strategy following  (Gupta et al. 2019). For each dataset $d$, let $f_{d}$ be the frequency of dataset $d$ in the mixed datasets. We define the dataset-level repeat factor as $r_{d}=\max\left(1,\sqrt{t/f_{d}}\right)$, where $t$ is a hyperparameter controlling the oversampling ratio. We then repeat dataset $d$ with a repeat factor of $r_{d}$.
A.2 More Model Details
@@
A.3 More Training Details
@@
The hyper-parameters in the multi-stage training are shown in Tab. 11. A simplified illustration of the multi-stage training is shown in Fig. 4.
Tab. 12: Comparison of Generic Segmentation. We compare different methods on the generic segmentation benchmarks.
@@
OMG-LLaVA(Zhang et al. 2024c) ConvNeXt-XXL 53.8 - -
PSALM(Zhang et al. 2024d) Swin-B 55.9 45.7 66.6
@@
Tab. 13: Comparison of OV Segmentation. We compare different methods on the A150-OV segmentation benchmarks.
@@
Tab. 14: Comparison of Reasoning Segmentation. We compare X-SAM with other methods on the reasoning segmentation benchmark.
@@
Tab. 15: Comparison of Interactive Segmentation. We compare X-SAM with other methods on the interactive segmentation benchmark.
@@
A.4 More Evaluation Details
PSALM COCO-VGD Evaluation. PSALM (Zhang et al. 2024d) is a segmentation MLLM that supports generic segmentation, open-vocabulary segmentation, referring segmentation, interactive segmentation, and more. To evaluate its performance on our proposed COCO-VGD dataset, we follow the same evaluation process as for X-SAM. We randomly sample some instance annotations, as done with X-SAM, and then feed them to PSALM to obtain instance-level predictions. PSALM is trained on the COCO-Interactive dataset, which shares the same source as COCO-VGD but performs classification for each interactive visual prompt. As a result, the AP metric for VGD segmentation is poor because the instance-level predictions are of low quality. This may explain why PSALM lacks the capability for instance-level visual grounding segmentation.
X-SAM COCO-Interactive Evaluation. X-SAM is the first unified segmentation MLLM, capable of adapting to all image segmentation tasks, including interactive segmentation (Zhang et al. 2024d). To evaluate its performance on the COCO-Interactive dataset, we first filter the instance-level predictions using a threshold of 0.5. We then calculate the IoU score between each remaining prediction and the visual prompt mask. Finally, we select the instance prediction with the highest IoU score as the final interactive segmentation result.
A.5 More Experimental Results
Tab. 16: Comparison of Closed-set Segmentation. We compare X-SAM with other methods on the COCO-Panoptic benchmark.
@@ A.5 More Experimental Results
Tab. 17: Comparison of Image-level Benchmarks. We compare X-SAM with other methods on the image-level benchmarks, including MME (Fu et al. 2024), MMBench (Liu et al. 2024c), SEED-Bench (Li et al. 2024a), POPE (Li et al. 2023b), and AI2D (Kembhavi et al. 2016).
@@
Generic Segmentation. Tab. 13 presents the results of generic segmentation. Thanks to our segmentor design and fine-tuning, X-SAM can adapt to generic segmentation and achieves competitive performance on the COCO-Panoptic dataset.
@@
Open-Vocabulary Segmentation. Tab. 13 shows the results of open-vocabulary segmentation. Benefiting from the robust mask generation of SAM and our mixed fine-tuning strategy, X-SAM achieves the best performance on open-vocabulary segmentation tasks.
@@
Reasoning Segmentation. In Tab. 14, we present the results of reasoning segmentation on the reasoning segmentation benchmark. We report the performance of our method on both the validation set and the test set, following (Lai et al. 2024). X-SAM achieves the best gIoU metric on both the validation and test sets, even though it is not specifically designed for reasoning segmentation. While the cIoU metric is not the best, it remains comparable to state-of-the-art methods. As the number of samples in the validation and test sets is limited, the results on this benchmark may not be stable.
@@
Interactive Segmentation. Tab. 15 shows the results of the interactive segmentation. Since interactive segmentation shares similar data with VGD segmentation, we exclude the training data for interactive segmentation and perform a process similar to that used for VGD segmentation to obtain the interactive segmentation results. X-SAM achieves the best or second-best performance on interactive segmentation, even without being trained on the specific data.
@@
Closed-Set Segmentation. In Tab. 16, we present the results of segmentor fine-tuning on the closed-set COCO-Panoptic benchmark. To preserve the robust generalization ability of SAM in mask prediction, we fine-tune the SAM encoder for only 36 epochs. Compared with other methods, our approach achieves comparable performance to the SAM-L encoder with just 36 epochs of fine-tuning.
@@
Image-level Benchmarks. In Tab. 17, we present the results of image-level benchmarks, including MME (Fu et al. 2024), MMBench (Liu et al. 2024c), SEED-Bench (Li et al. 2024a), POPE (Li et al. 2023b), and AI2D (Kembhavi et al. 2016). When jointly co-training with segmentation and conversation datasets, X-SAM achieves the best performance compared to other segmentation MLLMs on these benchmarks. Compared to LISA (Lai et al. 2024), PixelLM (Ren et al. 2024), and GLaMM (Rasheed et al. 2024), our method achieves significant improvements, demonstrating its effectiveness—even outperforming the previous best, OMG-LLaVA (Zhang et al. 2024c). On the POPE benchmark, X-SAM even surpasses LLaVA-V1.5 (Liu et al. 2023b), which is designed specifically for image-level conversation.
A.6 More Ablation Studies
Dataset Balance Resampling. In Tab. 18, we ablate the oversampling ratio $t$ from 0 to 0.1 in dataset balance resampling. When $t=0$, no dataset is oversampled. Otherwise, datasets are oversampled with a repeat factor of $r_{d}=\max\left(1,\sqrt{t/f_{d}}\right)$. We find that the performance on some small datasets is sensitive to the oversampling ratio $t$, especially for the reasoning segmentation dataset, where the gIoU score improves from 44.1% to 56.6% as $t$ increases from 0 to 0.1. Meanwhile, larger datasets are not sensitive to the oversampling ratio $t$. As a result, the overall performance improvement reaches its highest when $t=0.1$. Therefore, we set $t$ to 0.1 in the final experiment.
Image Encoder. In Tab. 19, we ablate the image encoder of X-SAM by replacing it with CLIP (Radford et al. 2021), SigLIP-so400m (Zhai et al. 2023), and SigLIP2-so400m (Tschannen et al. 2025). It can be observed that using more powerful image encoders improves the image content understanding ability of X-SAM, especially on the image conversation and GCG segmentation benchmarks, and even enhances performance on the generic segmentation benchmark. Although SigLIP achieves the best performance on the reasoning segmentation benchmark, it does not have a performance advantage on the other benchmarks. Meanwhile, SigLIP2 demonstrates more robust and consistently better performance across all benchmarks. Therefore, we adopt SigLIP2-so400m as the image encoder for the final experiment.
Tab. 18: Ablation of Dataset Balance Resampling. $\sum\Delta$ represents the total performance improvement compared to the first row.
@@ A.6 More Ablation Studies
PQ AP cIoU gIoU mIoU AP Acc
@@
Tab. 19: Ablation of Image Encoder. We ablate the impact of different image encoders, including CLIP (Radford et al. 2021), SigLIP (Zhai et al. 2023), and SigLIP2 (Tschannen et al. 2025), on the performance of X-SAM.
Img. Enc. COCO-Pan A150-OV RefCOCO Reason-Val GCG-Val COCO-VGD Conv.-MMB Conv.-MMB
PQ AP cIoU gIoU mIoU AP Acc Acc
@@
A.7 More Visualization Results
@@
Generic Segmentation. Fig. 5 shows the visualization results of X-SAM in generic segmentation, including semantic, instance and panoptic segmentation, which needs both the semantic and instance level understanding of the image. X-SAM can generate accurate and complete masks for the objects in the image.
@@
Open-Vocabulary Segmentation. Fig. 6 shows the visualization results of X-SAM in open-vocabulary (OV) segmentation, including OV-semantic, OV-instance, and OV-panoptic segmentation, which require segmenting objects that may not exist in the training set. X-SAM can segment objects that are not in the training set, demonstrating the robust generalization ability of the proposed method.
@@
GCG Segmentation. Fig. 7 shows the visualization results of X-SAM in GCG segmentation, which needs to describe the image and output the corresponding mask. X-SAM can not only effectively understand the image and generate the language description, but also generate the segmentation masks for the corresponding phrases.
@@
Referring Segmentation. Fig. 8 shows the visualization results of X-SAM in referring segmentation, which needs to segment objects that are referred to by natural language. X-SAM can effectively understand the referring expression and segment the objects that are referred to by natural language.
@@
Reasoning Segmentation. Fig. 9 shows the visualization results of X-SAM in reasoning segmentation, which needs to segment the object that is related to the question. X-SAM can effectively understand the complex question and then generate the corresponding mask for the question.
@@
Interactive Segmentation. Fig. 10 shows the visualization results of X-SAM in interactive segmentation, which needs to segment the individual objects with which the user interacts. X-SAM can generate the corresponding mask for the user’s interactive visual prompt.
@@
VGD Segmentation. Fig. 11 shows the visualization results of X-SAM in VGD segmentation for a single image, which needs to segment all the objects in the single image that is grounded with the user’s visual prompt. X-SAM can effectively segment all the objects in the image given by the user’s visual grounded prompt. In addition, X-SAM can perform VGD segmentation on the cross image, which needs to segment the objects grounded in another image. Fig. 12 shows the visualization results of X-SAM in VGD segmentation for cross image, which demonstrates the effectiveness of X-SAM for VGD segmentation both in single image and cross image.
A.8 Further Discussions
@@
Limitation. Although X-SAM achieves unified segmentation by extending segment anything to any segmentation, there is still considerable room for improvement. First, joint co-training with segmentation datasets and conversation datasets negatively impacts performance on some segmentation datasets, a phenomenon also observed in (Zhang et al. 2024c) and (Rasheed et al. 2024). This challenge may be addressed by designing a more balanced dataset mixture. Second, the performance of X-SAM is not optimal across all tasks, which has also been observed in other unified segmentation methods (Zhang et al. 2024c; Rasheed et al. 2024; Lai et al. 2024). This challenge remains a major obstacle for unified models and may be addressed by scaling up the model size and the amount of training data.

### v2

@@ Introduction
Mixed Fine-tuning. We ablate the impact of mixed fine-tuning on X-SAM’s performance. As shown in Table 6, mixed fine-tuning improves performance on out-of-domain COCO benchmarks, demonstrating X-SAM’s robust segmentation capabilities—for example, a 6.0% PQ increase on A150-OV and an 8.9% gIoU increase on Reason-Val. However, it results in a 0.8% PQ decrease on COCO-Pan due to the challenge of balancing performance in multi-sources training.
Dual Encoders. We ablate the design of the dual encoders in X-SAM. As shown in Table 7, dual encoders with either a SAM or Swin encoder benefit VGD segmentation, achieving 7.2% AP and 7.9% AP on COCO-VGD, respectively. Moreover, dual encoders with a SAM encoder consistently improve performance on GCG-Val and A150-OV, while the Swin encoder, which lacks robust segmentation capabilities, provides only a small improvement on A150-OV and even has a negative impact on GCG-Val.
Multi-stage Training. We ablate the impact of the multi-stage training strategy. As shown in Table 8, the S1 segmentor fine-tuning phase boosts the segmentation capability, producing a notable improvement of 9.3% PQ in COCO-Pan and 1.5% PQ in the A150-OV datasets. Meanwhile, the S2 alignment pre-training phase enhances image understanding capabilities, contributing an additional 2.1% Accuracy on Conv.-MMB. By integrating these stages, X-SAM demonstrates robust advances in image segmentation and comprehension, establishing its effectiveness in addressing complex visual tasks.
Segmentor Architecture. We ablate the impact of segmentor architecture by performing segmentor fine-tuning for 12 epochs. As shown in Table 9, M2F decoder brings a large improvement with 9.2% PQ as the effective design of M2F. The convolution connector performs better than the MLP connector, as the convolution spatial-awareness benefits segmentation, and multi-scale further improves the performance(10.7% PQ) with more diverse scale features.
@@
Table 8: Ablation on Multi-Stage(M-Stage) Training. S1: Stage 1, S2: Stage 2, S3: Stage 3, Conv.: conversation.
@@
Table 9: Ablation on Segmentor Architecture. Conn.: connector, M-Scale: multi-scale, Con.: convolution, M2F: Mask2Former.
Conclusion
@@
Acknowledgments
This work is supported by National Key Research and Development Program of China (2024YFE0203100), Scientific Research Innovation Capability Support Project for Young Faculty (No.ZYGXQNJSKYCXNLZCXM-I28), National Natural Science Foundation of China (NSFC) under Grants No.62476293, General Embodied AI Center of Sun Yat-sen University, National Natural Science Foundation of China (62402252) and (62536003) Guangdong High-Level Talent Programme (2024TQ08X283).
@@ Appendix A Technical Appendices and Supplementary Material
More Dataset Details
Training Datasets. In Table 10, we show the datasets used in our multi-stage training. For the segmentor fine-tuning stage, we fine-tune the segmentor on generic segmentation datasets. For the alignment pre-training stage, we pre-train the dual projectors on the alignment LLaVA 558K (Liu et al. 2023a) dataset. For the mixed fine-tuning stage, we fine-tune the whole model on mixed datasets, including both segmentation and conversation datasets. There are six types of datasets in total, including one image-level dataset and five segmentation datasets.
Building COCO-VGD Dataset. The COCO-VGD dataset is built on the images and annotations of the COCO2017 instance segmentation dataset, which provides instance-level segmentation masks for each object in an image. We automatically generate four types of visual prompts: point, scribe, box, and mask, for each instance in the image, following  (Zhang et al. 2024c). We randomly sample one type of visual prompt for each category as the visually grounded prompt during training and evaluation.
Dataset Balance Resampling. As shown in Table 10, the dataset sizes vary, ranging from 0.2K to 665K. The data ratio among the datasets is crucial to the model’s performance during mixed fine-tuning. To balance the different dataset sizes, we propose a dataset balance resampling strategy following  (Gupta et al. 2019). For each dataset $d$, let $f_{d}$ be the frequency of dataset $d$ in the mixed datasets. We define the dataset-level repeat factor as $r_{d}=\max\left(1,\sqrt{t/f_{d}}\right)$, where $t$ is a hyperparameter controlling the oversampling ratio. We then repeat dataset $d$ with a repeat factor of $r_{d}$.
More Model Details
@@
More Training Details
@@
The hyper-parameters in the multi-stage training are shown in Table 11. A simplified illustration of the multi-stage training is shown in Figure 5.
@@
OMG-LLaVA(Zhang et al. 2024b) ConvNeXt-XXL 53.8 - -
PSALM(Zhang et al. 2024c) Swin-B 55.9 45.7 66.6
@@
Table 12: Comparison of Generic Segmentation. We compare different methods on the generic segmentation benchmarks.
@@
Table 13: Comparison of OV Segmentation. We compare different methods on the A150-OV segmentation benchmarks.
@@
Table 14: Comparison of Reasoning Segmentation. We compare X-SAM with other methods on the reasoning segmentation benchmark.
@@
Table 15: Comparison of Interactive Segmentation. We compare X-SAM with other methods on the interactive segmentation benchmark.
More Evaluation Details
PSALM COCO-VGD Evaluation. PSALM (Zhang et al. 2024c) is a segmentation MLLM that supports generic segmentation, open-vocabulary segmentation, referring segmentation, interactive segmentation, and more. To evaluate its performance on our proposed COCO-VGD dataset, we follow the same evaluation process as for X-SAM. We randomly sample some instance annotations, as done with X-SAM, and then feed them to PSALM to obtain instance-level predictions. PSALM is trained on the COCO-Interactive dataset, which shares the same source as COCO-VGD but performs classification for each interactive visual prompt. As a result, the AP metric for VGD segmentation is poor because the instance-level predictions are of low quality. This may explain why PSALM lacks the capability for instance-level visual grounding segmentation.
X-SAM COCO-Interactive Evaluation. X-SAM is the first unified segmentation MLLM, capable of adapting to all image segmentation tasks, including interactive segmentation (Zhang et al. 2024c). To evaluate its performance on the COCO-Interactive dataset, we first filter the instance-level predictions using a threshold of 0.5. We then calculate the IoU score between each remaining prediction and the visual prompt mask. Finally, we select the instance prediction with the highest IoU score as the final interactive segmentation result.
More Experimental Results
LLM Zero-Shot val testA testB
cIoU gIoU cIoU gIoU cIoU gIoU
MattNet (Yu et al. 2018) - 47.5 48.2 58.7 59.3 45.3 46.1
LTS (Ding et al. 2021) - 52.3 52.7 61.9 62.6 49.9 50.4
CRIS (Wang et al. 2022) - 55.3 - 63.8 - 51.0 -
LISA (Lai et al. 2024) Vicuna-7B 38.7 32.2 52.6 48.5 44.8 39.7
X-SAM(Ours) Phi3-3.8B 36.9 44.8 32.3 34.5 36.4 38.9
Table 16: Comparison of Generalized Referring Segmentation. We compare X-SAM with other methods on the gRefCOCO benchmark.
@@
Table 17: Comparison of Closed-set Segmentation. We compare X-SAM with other methods on the COCO-Panoptic benchmark.
@@
Table 18: Comparison of Image-level Benchmarks. We compare X-SAM with other methods on the image-level benchmarks, including MME (Fu et al. 2024), MMBench (Liu et al. 2024b), SEED-Bench (Li et al. 2024b), POPE (Li et al. 2023), and AI2D (Kembhavi et al. 2016).
@@
Generic Segmentation. Table 13 presents the results of generic segmentation. Thanks to our segmentor design and fine-tuning, X-SAM can adapt to generic segmentation and achieves competitive performance on the COCO-Panoptic dataset.
@@
Open-Vocabulary Segmentation. Table 13 shows the results of open-vocabulary segmentation. Benefiting from the robust mask generation of SAM and our mixed fine-tuning strategy, X-SAM achieves the best performance on open-vocabulary segmentation tasks.
@@
Reasoning Segmentation. In Table 14, we present the results of reasoning segmentation on the reasoning segmentation benchmark. We report the performance of our method on both the validation set and the test set, following (Lai et al. 2024). X-SAM achieves the best gIoU metric on both the validation and test sets, even though it is not specifically designed for reasoning segmentation. While the cIoU metric is not the best, it remains comparable to state-of-the-art methods. As the number of samples in the validation and test sets is limited, the results on this benchmark may not be stable.
Generalized Referring Segmentation. In Table 16, we present the results of generalized referring segmentation on the gRefCOCO benchmark. We report the performance of our method on both the validation set and the test set, following (Yu et al. 2018). X-SAM achieves the comparable performance to state-of-the-art methods under zero-shot setting, it will achieve better performance with the fine-tuning on the gRefCOCO dataset.
@@
Interactive Segmentation. Table 15 shows the results of the interactive segmentation. Since interactive segmentation shares similar data with VGD segmentation, we exclude the training data for interactive segmentation and perform a process similar to that used for VGD segmentation to obtain the interactive segmentation results. X-SAM achieves the best or second-best performance on interactive segmentation, even without being trained on the specific data.
@@
Closed-Set Segmentation. In Table 17, we present the results of segmentor fine-tuning on the closed-set COCO-Panoptic benchmark. To preserve the robust generalization ability of SAM in mask prediction, we fine-tune the SAM encoder for only 36 epochs. Compared with other methods, our approach achieves comparable performance to the SAM-L encoder with just 36 epochs of fine-tuning without any complex data augmentation.
@@
Image-level Benchmarks. In Table 18, we present the results of image-level benchmarks, including MME (Fu et al. 2024), MMBench (Liu et al. 2024b), SEED-Bench (Li et al. 2024b), POPE (Li et al. 2023), and AI2D (Kembhavi et al. 2016). When jointly co-training with segmentation and conversation datasets, X-SAM achieves the best performance compared to other segmentation MLLMs on these benchmarks. Compared to LISA (Lai et al. 2024), PixelLM (Ren et al. 2024), and GLaMM (Rasheed et al. 2024), our method achieves significant improvements, demonstrating its effectiveness—even outperforming the previous best, OMG-LLaVA (Zhang et al. 2024b). On the POPE benchmark, X-SAM even surpasses LLaVA-V1.5 (Liu et al. 2023a), which is designed specifically for image-level conversation.
@@
PQ PQ cIoU gIoU mIoU AP Acc
@@
Table 19: Ablation of Dataset Balance Resampling. $\sum\Delta$ represents the total performance improvement compared to the first row.
Img. Enc. COCO-Pan A150-OV RefCOCO Reason-Val GCG-Val COCO-VGD Conv.-MME Conv.-MMB
PQ PQ cIoU gIoU mIoU AP Acc Acc
@@
Table 20: Ablation of Image Encoder. We ablate the impact of different image encoders, including CLIP (Radford et al. 2021), SigLIP (Zhai et al. 2023), and SigLIP2 (Tschannen et al. 2025), on the performance of X-SAM.
LLM COCO-Pan COCO-VGD Conv.-MMB
PQ AP Acc
Phi3-3.8B (Abdin et al. 2024) 54.7 47.9 69.3
Qwen3-1.7B (Bai et al. 2023) 54.3 47.8 65.8
Vicuna-7B (Chiang et al. 2023) 54.5 47.3 71.3
Table 21: Ablation of LLMs. We ablate the performance of X-SAM with different LLMs.
Seg Data Conv.-MME SEED POPE AI2D
Acc Acc Acc Acc
1397 / 332 69.0 87.6 64.8
1374 / 312 69.3 89.3 62.6
Table 22: Ablation of Segmentation Data. We ablate the performance of X-SAM with and without segmentation data on the image-level benchmarks.
More Ablation Studies
Dataset Balance Resampling. In Table 19, we ablate the oversampling ratio $t$ from 0 to 0.1 in dataset balance resampling. When $t=0$, no dataset is oversampled. Otherwise, datasets are oversampled with a repeat factor of $r_{d}=\max\left(1,\sqrt{t/f_{d}}\right)$. We find that the performance on some small datasets is sensitive to the oversampling ratio $t$, especially for the reasoning segmentation dataset, where the gIoU score improves from 44.1% to 56.6% as $t$ increases from 0 to 0.1. Meanwhile, larger datasets are not sensitive to the oversampling ratio $t$. As a result, the overall performance improvement reaches its highest when $t=0.1$. Therefore, we set $t$ to 0.1 in the final experiment.
Image Encoder. In Table 20, we ablate the image encoder of X-SAM by replacing it with CLIP (Radford et al. 2021), SigLIP-so400m (Zhai et al. 2023), and SigLIP2-so400m (Tschannen et al. 2025). It can be observed that using more powerful image encoders improves the image content understanding ability of X-SAM, especially on the image conversation and GCG segmentation benchmarks, and even enhances performance on the generic segmentation benchmark. Although SigLIP achieves the best performance on the reasoning segmentation benchmark, it does not have a performance advantage on the other benchmarks. Meanwhile, SigLIP2 demonstrates more robust and consistently better performance across all benchmarks. Therefore, we adopt SigLIP2-so400m as the image encoder for the final experiment.
LLM. In Table 22, we ablate the LLM of X-SAM with different LLMs. We find that Phi3-3.8B (Abdin et al. 2024) achieves the best performance on the COCO-Pan and COCO-VGD benchmarks, while Vicuna-7B (Chiang et al. 2023) achieves the best performance on the Conv.-MMB benchmark. Moreover, Phi3-3.8B show more robust and consistent performance across segmentation benchmarks, we adopt it as the LLM for the final experiment.
Segmentation Data. In Table 22, we ablate the impact of segmentation data on the image-level benchmarks. We find that adding segmentation data improves the performance of X-SAM on the image-level benchmarks, especially on the hallucination-prone benchmarks, such as SEED and POPE.
More Visualization Results
@@
Generic Segmentation. Figure 6 shows the visualization results of X-SAM in generic segmentation, including semantic, instance and panoptic segmentation, which needs both the semantic and instance level understanding of the image. X-SAM can generate accurate and complete masks for the objects in the image.
@@
Open-Vocabulary Segmentation. Figure 7 shows the visualization results of X-SAM in open-vocabulary (OV) segmentation, including OV-semantic, OV-instance, and OV-panoptic segmentation, which require segmenting objects that may not exist in the training set. X-SAM can segment objects that are not in the training set, demonstrating the robust generalization ability of the proposed method.
@@
GCG Segmentation. Figure 8 shows the visualization results of X-SAM in GCG segmentation, which needs to describe the image and output the corresponding mask. X-SAM can not only effectively understand the image and generate the language description, but also generate the segmentation masks for the corresponding phrases.
@@
Referring Segmentation. Figure 9 shows the visualization results of X-SAM in referring segmentation, which needs to segment objects that are referred to by natural language. X-SAM can effectively understand the referring expression and segment the objects that are referred to by natural language.
@@
Reasoning Segmentation. Figure 10 shows the visualization results of X-SAM in reasoning segmentation, which needs to segment the object that is related to the question. X-SAM can effectively understand the complex question and then generate the corresponding mask for the question.
@@
Interactive Segmentation. Figure 11 shows the visualization results of X-SAM in interactive segmentation, which needs to segment the individual objects with which the user interacts. X-SAM can generate the corresponding mask for the user’s interactive visual prompt.
@@
VGD Segmentation. Figure 12 shows the visualization results of X-SAM in VGD segmentation for a single image, which needs to segment all the objects in the single image that is grounded with the user’s visual prompt. X-SAM can effectively segment all the objects in the image given by the user’s visual grounded prompt. In addition, X-SAM can perform VGD segmentation on the cross image, which needs to segment the objects grounded in another image. Figure 13 shows the visualization results of X-SAM in VGD segmentation for cross image, which demonstrates the effectiveness of X-SAM for VGD segmentation both in single image and cross image.
Further Discussions
@@
Limitation. Although X-SAM achieves unified segmentation by extending segment anything to any segmentation, there is still considerable room for improvement. First, joint co-training with segmentation datasets and conversation datasets negatively impacts performance on some segmentation datasets, a phenomenon also observed in (Zhang et al. 2024b) and (Rasheed et al. 2024). This challenge may be addressed by designing a more balanced dataset mixture. Second, the performance of X-SAM is not optimal across all tasks, which has also been observed in other unified segmentation methods (Zhang et al. 2024b; Rasheed et al. 2024; Lai et al. 2024). This challenge remains a major obstacle for unified models and may be addressed by scaling up the model size and the amount of training data.